HETAV PATEL

Grade 11,

Compute CM of two equal masses 6kg weight separate by a distance of 1.5m. where is the Center of Mass. Where is the Centre of Mass if masses are 5kg and 8kg.

Compute CM of two equal masses 6kg weight separate by a distance of 1.5m. where is the Center of  Mass.
Where is the Centre of Mass if masses are 5kg and 8kg.

Grade:12th pass

1 Answers

Aman Patel
5 years ago
Just u need to apply formula...In the first case centre of mass will be at 1.5/2m....and in second case =5(0)+8(1.5)/5+8=12/13